OpenStudy (anonymous):

It was, what is the derivative of cos^2(x)^3?

OpenStudy (anonymous):

so its like cos ^ 6 x ? thats 6 cos x ^ 5 (-sin x)

OpenStudy (anonymous):

do you mean \[\cos^2(x^3)\] ???

OpenStudy (anonymous):

\[= -6x^2 \cos(x^3)\sin(x^3)\]